THE MOLECULAR FRAME ELECTRIC DIPOLE MOMENT AND HYPERFINE INTERACTIONS IN HAFNIUM FLUORIDE, HfF

\underline{ANH LE} AND TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; LEONID SKRIPNIKOV AND ANATOLY V. TITOV, Petersburg Nuclear Physics Institute, Gatchina, 188300, Russia and Quantum Mechanics Division, St. Petersburg State University, St. Petersburg 198904, Russia.

The identification of HfF^{+} as a possible candidate for a d_{e} measurement has stimulated new interest in the spectroscopy of both HfF^{+} ^{a}^{,} ^{b}^{,} ^{c} and neutral HfF ^{a,} ^{d}. Studies of the neutral are relevant because photoionization schemes can be used to produce the cations. More importantly, computational methodologies used to predict the electronic wavefunction of HfF^{+} can be effectively assessed by making a comparison of predicted and experimental properties of the neutral, which are more readily determinable. The (1,0)[17.9]2.5 -X^{2}Δ_{3/2} band of hafnium monofluoride (HfF) has been recorded using high-resolution laser-induced fluorescence spectroscopy both field-free and in the presence of a static electric field. The field-free spectra of ^{177}HfF, ^{179}HfF, and ^{180}HfF were model to generate a set of fine and hyperfine parameters for the X^{2}Δ_{3/2} (v=0) and [17.9]2.5 (v=1) states. The observed optical Stark shifts for the ^{180}HfF isotopologue were analyzed to produce the molecular frame electric dipole moments of 1.66(1)D and 0.419(7)D for the X^{2}Δ_{3/2} and [17.9]2.5 states, respectively. A two-step ab initio calculation consisting of a two-component generalized relativistic effective core potential calculation (GRECP) followed by a restoration of the proper four-component wavefunction was performed to predict the properties of ground state HfF.

aB. J. Barker, I. O. Antonov, V. E. Bondybey, and M. C. Heaven, J. Chem. Phys., 134, 201102 (2011).

bK. C. Cossel, D. N. Gresh, L. C. Sinclair, T. Coffey, L. V. Skripnikov, A. N. Petrov, N. S. Mosyagin, A. V. Titov, R. W. Field, E. R. Meyer, E. A. Cornell and J. Ye, Chem. Phys. Lett., 546, 1 (2012).

cM. Grau, A. E. Leanhardt, H. Loh, L. C. Sinclair, R. P. Stutz, T. S. Yahn, and E. A. Cornell, J. Mol. Spectroc., 272, 32 (2012).

dH. Loh, R. P. Stutz, T. S. Yahn, H. Looser, R. W. Field, and E. A. Cornell, J. Mol. Spectroc.,276-277, 49 (2012).